# Finitistic dimensions over commutative DG-rings

### v3

@@ 6. Constructing DG-modules of prescribed projective dimension
We warn the reader that unlike in the case of rings, in general local-Cohen-Macaulay does not imply Cohen-Macaulay due to localization causing a drop in amplitude, see [21, Example 8.3]. However, when the spectrum of $\mathrm{H}^{0}(A)$ is irreducible and $\mathrm{H}^{0}(A)$ is catenary, the two definitions agree by combining [21, Corollary 8.7] and [23, Theorem 2.5].
@@
Building on recent results on the loci of Cohen-Macaulay DG-rings [24], the following lemma provides a DG-version of [3, Proposition 5.2].
@@
By localizing $A$ at a height $n$ prime of $\mathrm{H}^{0}(A)$, we may assume $(A,\bar{\mathfrak{m}})$ is local and $\dim(\mathrm{H}^{0}(A))=n$. We write $\widehat{A}:=\mathsf{L}\Lambda(A,\bar{\mathfrak{m}})$ for the derived $\bar{\mathfrak{m}}$-completion of $A$. Note that  [20, Proposition 4.16] tells us that $\mathrm{H}^{0}(\widehat{A})=\widehat{\mathrm{H}^{0}(A)}$ and $\mathrm{H}^{0}(A)\to\mathrm{H}^{0}(\widehat{A})$ is the usual $\mathfrak{m}$-adic completion map of rings. To obtain the prime $\bar{\mathfrak{p}}$ as in the statement we use induction to construct a strict chain of length $n$
@@
in $\widehat{\mathrm{H}^{0}(A)}$ which lies over the given chain in $\mathrm{H}^{0}(A)$. Clearly, $\bar{\mathrm{Q}}_{0}\in\mathrm{LCM}(\widehat{A})\cap\mathrm{CM}(\mathrm{H}^{0}(\widehat{A}))$ and $\bar{\mathrm{P}}_{0}=\bar{\mathrm{Q}}_{0}\cap A$ has height zero. Set $\bar{\mathfrak{q}}_{0}:=\bar{\mathrm{Q}}_{0}$. By [19, Proposition 7.21] the DG-ring $\widehat{A}$ has a dualizing DG-module, so we may apply [24, Theorem 12] to choose an open neighbourhood $D(f)$ of $\bar{\mathfrak{q}}_{0}$ in $\mathrm{LCM}(\widehat{A})\cap\mathrm{CM}(\mathrm{H}^{0}(\widehat{A}))$ which completes the base case.
@@
$\mathsf{X}_{w}=\bigcup_{\mathrm{ht}(\bar{\mathfrak{p}})=w}\{\bar{\mathfrak{p}}\in\mathrm{Spec}(\mathrm{H}^{0}(A))\mid\bar{\mathfrak{q}}_{w-1}\subseteq\bar{\mathfrak{p}}\subseteq\bar{\mathrm{Q}}_{w+1}\text{ and }f\in\bar{\mathfrak{p}}\}.$
@@
As $A\rightarrow\widehat{A}$ is flat by [18, Corollary 4.6], we can now use 6.1 to conclude that $\bar{\mathfrak{q}}_{i}\cap\mathrm{H}^{0}(A)\in\mathrm{LCM}(A)\cap\mathrm{CM}(\mathrm{H}^{0}(A))$, which is what we wanted to show. ∎
@@
The following classical fact was proved in [3, Proposition 5.4].
@@
If $n=0$, $M=A$ will do. So assume $n\geq 1$. Choose a prime $\bar{\mathfrak{p}}\subseteq\mathrm{H}^{0}(A)$ as in 6.2. By [21, Corollary 5.21], there exists a sequence $\underline{a}=a_{1},\dots,a_{n-1}\in\mathrm{H}^{0}(A)$ whose image in $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$ forms an $A_{\bar{\mathfrak{p}}}$ regular sequence while simultaneously being a system of parameters for $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$. Since $\bar{\mathfrak{p}}\in\mathrm{LCM}(\mathrm{H}^{0}(A))$, the image of this sequence in $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$ is also $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$-regular.
@@
$\operatorname{proj\,dim}_{A}(K(A_{s};a_{1},\dots,a_{n-1})=\operatorname{proj\,dim}_{\mathrm{H}^{0}(A)}(K(\mathrm{H}^{0}(A)_{xf};a_{1},\dots,a_{n-1})))=n$
@@
Since $\mathrm{H}^{0}(A)\otimes^{\mathsf{L}}_{A}M\in\mathsf{D}_{\mathrm{f}}^{\mathrm{-}}(\mathrm{H}^{0}(A))$ by 3.1, by applying [1, Corollary 2.10.F] we have
@@ 7.1. Finitistic projective dimensions
Let us contrast the above bounds to those in the case of modules, or even complexes, over commutative noetherian rings. We highlight that the bounds in the above theorem cannot be improved, as is shown in Section 8. In particular, we may find commutative noetherian DG-rings with bounded cohomology whose big finitistic projective dimension is strictly less than $\mathrm{dim}(\mathrm{H}^{0}(A))$. This is a stark difference to the case of commutative noetherian rings. If $A$ is an ordinary commutative noetherian ring, then it is known (see, for example, [5, E8.5.4]) that the finitistic projective dimension of $A$ (defined as in 7.1) is equal to $\dim(A)$. Note that this is incorporated into the preceding theorem, simply by considering the case when the DG-ring has amplitude zero.
We now turn to identifying the small finitistic projective dimension. To begin with, we treat the local case. The following is an easy consequence of the DG-Auslander-Buchsbaum formula [14, Theorem 2.16].
@@
By combining the previous result with [21, Corollary 5.5] we obtain the following characterization of local-Cohen-Macaulay DG-rings.
@@
By applying 2.3 and [1, Proposition 5.3.P], we have
@@ 7.2. Finitistic flat and injective dimensions
For brevity we write $\mathbb{I}M=\mathsf{R}\mathrm{Hom}_{A}(M,E)$. We refer the reader to [4, Section 3.1] and [19] for more details on this construction. It follows from (ii) that we have the equalities
@@
The equality $\mathsf{fpd}(A)=\mathsf{ffd}(A)$ is 6.6, which are equal to $\operatorname{seq.depth}_{A}(A)-\mathrm{amp}(A)$ by 7.5. For the injective case, let $M\in\mathsf{D}_{\mathrm{f}}^{\mathrm{b}}(A)$ have finite injective dimension. By the Bass formula given in [14, Theorem 3.33], we have that $\operatorname{inj\,dim}_{A}(M)-\sup(M)=\operatorname{seq.depth}_{A}(A)-\operatorname{amp}(A)$. Therefore by the definition of $\mathsf{fid}(A)$, it is clear that $\mathsf{fid}(A)=\operatorname{seq.depth}_{A}(A)-\text{amp}(A)$. ∎
@@
We also have the following results for non-local DG-rings, whose proofs are analogous to that of 7.8, using [1, Propositions 5.3.F and 5.3.I].
@@ 8. Proving optimality of the bounds
Take $A$ to be the trivial extension DG-ring $A=R\ltimes\!C[n]$; see [11, Section 1] for details. Note that $A$ is a commutative noetherian DG-ring with $\mathrm{H}^{0}(A)=R$, $\mathrm{H}^{-n}(A)=C$, and $\mathrm{H}^{i}(A)=0$ for all $i\notin\{-n,0\}$. In particular, $\operatorname{amp}(A)=n$ and $\dim(\mathrm{H}^{0}(A))=d$.
@@
Since $A$ and $\mathrm{H}^{0}(A)$ are Gorenstein, it follows that $A$ is a dualizing DG-module over itself and that $\mathrm{H}^{0}(A)$ is dualizing complex over itself. By [9, Theorem I] or [26, Proposition 7.5] and the fact that over a local ring a dualizing complex is unique up to a shift, it follows that $\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),A)\cong\mathrm{H}^{0}(A)[k]$ for some $k\in\mathbb{Z}$. By 3.1 we know that
@@
To begin with, consider a Gorenstein local ring $B$ of dimension $d+k$ for some $k>0$ and let $b_{1},\dots,b_{k}$ be a regular sequence in $B$ generating an ideal $I$. The quotient ring $B/I$ is also Gorenstein, and has dimension $d$. If we enlarge the generating set $b_{1},\dots,b_{k}$ to a generating set $b_{1},\dots,b_{n+k}$ for $I$ and set $A:=K(B;b_{1},\dots,b_{n+k})$, then $A$ is a commutative noetherian local DG-ring with bounded cohomology satisfying $\dim(\mathrm{H}^{0}(A))=d$ and $\operatorname{amp}(A)=n$. By [10, Theorem 4.9] or [2], the fact that $B$ is a Gorenstein ring implies that $A$ is a Gorenstein DG-ring. Moreover, $\mathrm{H}^{0}(A)=B/I$ is also Gorenstein, and hence $\mathsf{FPD}(A)=\dim(\mathrm{H}^{0}(A))-\operatorname{amp}(A)$ by 8.3.

### v4

@@ 6. Constructing DG-modules of prescribed projective dimension
We warn the reader that unlike in the case of rings, in general local-Cohen-Macaulay does not imply Cohen-Macaulay due to localization causing a drop in amplitude, see [22, Example 8.3]. However, when the spectrum of $\mathrm{H}^{0}(A)$ is irreducible and $\mathrm{H}^{0}(A)$ is catenary, the two definitions agree by combining [22, Corollary 8.7] and [25, Theorem 2.5].
@@
Building on recent results on the loci of Cohen-Macaulay DG-rings [24], the following lemma provides a DG-version of [4, Proposition 5.2].
@@
By localizing $A$ at a height $n$ prime of $\mathrm{H}^{0}(A)$, we may assume $(A,\bar{\mathfrak{m}})$ is local and $\dim(\mathrm{H}^{0}(A))=n$. We write $\widehat{A}:=\mathsf{L}\Lambda(A,\bar{\mathfrak{m}})$ for the derived $\bar{\mathfrak{m}}$-completion of $A$. Note that  [21, Proposition 4.16] tells us that $\mathrm{H}^{0}(\widehat{A})=\widehat{\mathrm{H}^{0}(A)}$ and $\mathrm{H}^{0}(A)\to\mathrm{H}^{0}(\widehat{A})$ is the usual $\mathfrak{m}$-adic completion map of rings. To obtain the prime $\bar{\mathfrak{p}}$ as in the statement we use induction to construct a strict chain of length $n$
@@
in $\widehat{\mathrm{H}^{0}(A)}$ which lies over the given chain in $\mathrm{H}^{0}(A)$. Clearly, $\bar{\mathrm{Q}}_{0}\in\mathrm{LCM}(\widehat{A})\cap\mathrm{CM}(\mathrm{H}^{0}(\widehat{A}))$ and $\bar{\mathrm{P}}_{0}=\bar{\mathrm{Q}}_{0}\cap A$ has height zero. Set $\bar{\mathfrak{q}}_{0}:=\bar{\mathrm{Q}}_{0}$. By [20, Proposition 7.21] the DG-ring $\widehat{A}$ has a dualizing DG-module, so we may apply [24, Theorem 12] to choose an open neighbourhood $D(f)$ of $\bar{\mathfrak{q}}_{0}$ in $\mathrm{LCM}(\widehat{A})\cap\mathrm{CM}(\mathrm{H}^{0}(\widehat{A}))$ which completes the base case.
@@
$\mathsf{X}_{w}=\{\bar{\mathfrak{p}}\in\mathrm{Spec}(\mathrm{H}^{0}(A))\mid\bar{\mathfrak{q}}_{w-1}\subseteq\bar{\mathfrak{p}}\subseteq\bar{\mathrm{Q}}_{w+1},\mathrm{ht}(\bar{\mathfrak{p}})=w\text{ and }f\in\bar{\mathfrak{p}}\}.$
@@
As $A\rightarrow\widehat{A}$ is flat by [19, Corollary 4.6], we can now use 6.1 to conclude that $\bar{\mathfrak{q}}_{i}\cap\mathrm{H}^{0}(A)\in\mathrm{LCM}(A)\cap\mathrm{CM}(\mathrm{H}^{0}(A))$, which is what we wanted to show. ∎
@@
The following classical fact was proved in [4, Proposition 5.4].
@@
If $n=0$, $M=A$ will do. So assume $n\geq 1$. Choose a prime $\bar{\mathfrak{p}}\subseteq\mathrm{H}^{0}(A)$ as in 6.2. By [22, Corollary 5.21], there exists a sequence $\underline{a}=a_{1},\dots,a_{n-1}\in\mathrm{H}^{0}(A)$ whose image in $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$ forms an $A_{\bar{\mathfrak{p}}}$ regular sequence while simultaneously being a system of parameters for $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$. Since $\bar{\mathfrak{p}}\in\mathrm{LCM}(\mathrm{H}^{0}(A))$, the image of this sequence in $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$ is also $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$-regular.
@@
$\operatorname{proj\,dim}_{A}(K(A_{s};a_{1},\dots,a_{n-1}))=\operatorname{proj\,dim}_{\mathrm{H}^{0}(A)}(K(\mathrm{H}^{0}(A)_{xf};a_{1},\dots,a_{n-1}))=n$
@@
Since $\mathrm{H}^{0}(A)\otimes^{\mathsf{L}}_{A}M\in\mathsf{D}_{\mathrm{f}}^{\mathrm{-}}(\mathrm{H}^{0}(A))$ by 3.1, by applying [2, Corollary 2.10.F] we have
@@ 7.1. Finitistic projective dimensions
Let us contrast the above bounds to those in the case of modules, or even complexes, over commutative noetherian rings. We highlight that the bounds in the above theorem cannot be improved, as is shown in Section 8. In particular, we may find commutative noetherian DG-rings with bounded cohomology whose big finitistic projective dimension is strictly less than $\mathrm{dim}(\mathrm{H}^{0}(A))$. This is a stark difference to the case of commutative noetherian rings. If $A$ is an ordinary commutative noetherian ring, then it is known (see, for example, [6, E8.5.4]) that the finitistic projective dimension of $A$ (defined as in 7.1) is equal to $\dim(A)$. Note that this is incorporated into the preceding theorem, simply by considering the case when the DG-ring has amplitude zero.
We now turn to identifying the small finitistic projective dimension. To begin with, we treat the local case. The following is an easy consequence of the DG-Auslander-Buchsbaum formula [15, Theorem 2.16].
@@
By combining the previous result with [22, Corollary 5.5] we obtain the following characterization of local-Cohen-Macaulay DG-rings.
@@
By applying 2.3 and [2, Proposition 5.3.P], we have
@@ 7.2. Finitistic flat and injective dimensions
For brevity we write $\mathbb{I}M=\mathsf{R}\mathrm{Hom}_{A}(M,E)$. We refer the reader to [5, Section 3.1] and [20] for more details on this construction. It follows from (ii) that we have the equalities
@@
The equality $\mathsf{fpd}(A)=\mathsf{ffd}(A)$ is 6.6, which are equal to $\operatorname{seq.depth}_{A}(A)-\mathrm{amp}(A)$ by 7.5. For the injective case, let $M\in\mathsf{D}_{\mathrm{f}}^{\mathrm{b}}(A)$ have finite injective dimension. By the Bass formula given in [15, Theorem 3.33], we have that $\operatorname{inj\,dim}_{A}(M)-\sup(M)=\operatorname{seq.depth}_{A}(A)-\operatorname{amp}(A)$. Therefore by the definition of $\mathsf{fid}(A)$, it is clear that $\mathsf{fid}(A)=\operatorname{seq.depth}_{A}(A)-\text{amp}(A)$. ∎
@@
We also have the following results for non-local DG-rings, whose proofs are analogous to that of 7.8, using [2, Propositions 5.3.F and 5.3.I].
@@ 8. Proving optimality of the bounds
Take $A$ to be the trivial extension DG-ring $A=R\ltimes\!C[n]$; see [12, Section 1] for details. Note that $A$ is a commutative noetherian DG-ring with $\mathrm{H}^{0}(A)=R$, $\mathrm{H}^{-n}(A)=C$, and $\mathrm{H}^{i}(A)=0$ for all $i\notin\{-n,0\}$. In particular, $\operatorname{amp}(A)=n$ and $\dim(\mathrm{H}^{0}(A))=d$.
@@
Since $A$ and $\mathrm{H}^{0}(A)$ are Gorenstein, it follows that $A$ is a dualizing DG-module over itself and that $\mathrm{H}^{0}(A)$ is dualizing complex over itself. By [10, Theorem I] or [27, Proposition 7.5] and the fact that over a local ring a dualizing complex is unique up to a shift, it follows that $\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),A)\cong\mathrm{H}^{0}(A)[k]$ for some $k\in\mathbb{Z}$. By 3.1 we know that
@@
To begin with, consider a Gorenstein local ring $B$ of dimension $d+k$ for some $k>0$ and let $b_{1},\dots,b_{k}$ be a regular sequence in $B$ generating an ideal $I$. The quotient ring $B/I$ is also Gorenstein, and has dimension $d$. If we enlarge the generating set $b_{1},\dots,b_{k}$ to a generating set $b_{1},\dots,b_{n+k}$ for $I$ and set $A:=K(B;b_{1},\dots,b_{n+k})$, then $A$ is a commutative noetherian local DG-ring with bounded cohomology satisfying $\dim(\mathrm{H}^{0}(A))=d$ and $\operatorname{amp}(A)=n$. By [11, Theorem 4.9] or [3], the fact that $B$ is a Gorenstein ring implies that $A$ is a Gorenstein DG-ring. Moreover, $\mathrm{H}^{0}(A)=B/I$ is also Gorenstein, and hence $\mathsf{FPD}(A)=\dim(\mathrm{H}^{0}(A))-\operatorname{amp}(A)$ by 8.3.